In exponential form.
What is the cube root of X^2Y^2?

Answer by mukhopadhyay(490)

Cube root of x^2y^2
= (x^2*y^2)^(1/3)
= (x^2)^(1/3) * (y^2)^(1/3)
= x^(2/3)*y^(2/3)